# Knockoffs for low dimensions: changing the nominal level post-hoc to gain power while controlling the FDR

### v1

@@ 7 Discussion
In this paper, we have shown how the significance level of the knockoff method can be chosen data-dependently while maintaining a reasonable FDR control. Remarkably, this flexibility comes without any price: (1) if the original method does not make any rejections, we can increase the significance level to potentially make rejections and (2) if the original method makes rejections, we can make the same rejections while often reducing (and never enlarging) the significance level, thus increasing the precision. In this way, the results obtained by our post-hoc knockoff procedure are always more informative than the results offered by the original knockoff procedure. This distinguishes our method from existing post-hoc $\alpha$ approaches for other testing problems [8, 14, 35, 6, 5], where the gained flexibility usually reduces the power compared to the standard methods.
Our post-hoc approach solves the well-known issue of the original knockoff method, which requires at least $1/\alpha^{\textnormal{kn}}$ rejections to make any rejection. This makes our framework particularly suitable for settings with small numbers of variables and/or sparse signals.
Furthermore, we integrated the post-hoc idea into derandomized knockoff procedures. While the latter require to choose some arbitrary parameters ($\alpha_{\textnormal{ebh}}$ for FDR and $\eta$ for PFER control) that influences the power and the guarantee of the procedure, we showed that the parameters can be chosen completely data-dependent while maintaining a reasonable control. Ren et al. [22] even mentioned that a data-dependent choice of the parameter $\eta$ would be desirable but is not allowed and thus proposed data-splitting to determine $\eta$ in a data-driven way. Obviously, our approach to choose $\eta$ based on the entire data is much more powerful and practical.
In Section 4.1 we introduced a uniform improvement of the derandomized (ph-)knockoff methods based on the e-Closure Principle. However, due to its high computational requirements, it is difficult to apply in practice. We wonder whether there are computationally efficient short-cuts for this method. This could be explored in future work.

### v2

@@ 7 Discussion
In this paper, we have shown how the significance level of the knockoff method can be chosen data-dependently while maintaining a reasonable FDR control. Remarkably, this flexibility comes without any price: (1) if the original method does not make any rejections, we can increase the significance level to potentially make rejections and (2) if the original method makes rejections, we can make the same rejections while often reducing (and never enlarging) the significance level, thus increasing the precision. In this way, the results obtained by our post-hoc knockoff procedure are always more informative than the results offered by the original knockoff procedure. This distinguishes our method from existing post-hoc $\alpha$ approaches for other testing problems (Grünwald, 2024; Koning, 2023; Xu and Ramdas, 2022; Gauthier et al., 2025; Dhillon et al., 2025), where the gained flexibility usually reduces the power compared to the standard methods.
Our post-hoc approach solves the well-known issue of the original knockoff method, which requires at least $1/\alpha^{\textnormal{kn}}$ rejections to make any rejection. This makes our framework particularly suitable for settings with small numbers of variables and/or sparse signals. Furthermore, we integrated the post-hoc idea into derandomized knockoff procedures. While the latter require the choice of some arbitrary parameters ($\alpha_{\textnormal{ebh}}$ for FDR and $\eta$ for PFER control), which influences the power and the guarantee of the procedure, we showed that the parameters can be chosen in a fully data-dependent manner while maintaining reasonable control. Ren et al. (2023) even mentioned that a data-dependent choice of the parameter $\eta$ would be desirable but is not allowed and thus proposed data-splitting to determine $\eta$ in a data-driven way. Our approach to choose $\eta$ based on the entire data is much more powerful and practical.
In Section 4.2, we introduced a uniform improvement of the derandomized (ph‑)knockoff methods based on the e‑Closure Principle. However, due to its substantial computational demands, the approach remains challenging to apply in practice. Identifying computationally efficient shortcuts for this method is an interesting direction for future work.
Acknowledgments
We are thankful to Zhimei Ren for bringing the conditional calibration paper by Luo et al. (2025) to our attention, and we also thank Alain Marti for his valuable feedback on the manuscript.